# Locally Differentially Private Thresholding Bandits

### v1

@@ 2.2 Thresholding Bandits
In the fixed budget thresholding bandit problem (Locatelli et al., 2016), the goal is to identify a subset of arms from a set $\mathcal{A}=\{1,2,\dots,K\}$ whose mean rewards exceed a prespecified threshold, up to a tolerance factor $\zeta$. Formally, let each arm $i$ in $\mathcal{A}$ have a reward that is distributed within [0,1][0,1][ 0 , 1 ] and an unknown mean reward $\mu_{i}$. Let $\tau$ denote the desired threshold, and let $\Delta_{i}=|\tau-\mu_{i}|+\zeta$ represent the deviation of the mean reward of arm $i$ from the threshold plus some tolerance $\zeta\geq 0$. The objective is to return $\mathcal{S}_{\tau}\subseteq\mathcal{A}$ where:
@@ 2.3.1 Bernoulli Mechanism
The Bernoulli mechanism, which will be utilized in both the fixed budget (Section 3.1) and fixed confidence (Section 4.1) settings, transforms a bounded reward in [0,1][0,1][ 0 , 1 ] into a Bernoulli random variable that satisfies differential privacy guarantees. This mechanism produces a binary output (0 or 1) that serves as a privacy-preserving approximation of the reward. Consequently, the agent does not have direct access to the true reward of the arm; instead, they observe only the realization of a Bernoulli random variable.
@@
[…]
@@
[…]
@@ 3.1 Algorithm for the Fixed Budget Case
where $T_{k}(t-1)$ is the number of pulls of arm $k$ up to time $t-1$ and
@@
is the estimate of arm $k$’s private reward deviation from $\tau_{\epsilon}$ at time $t-1$, and
@@
6:  Pull arm $I_{t}\leftarrow\arg\min_{k\leq K}B_{k}(t)$triangleright where $B_{k}(t)\leftarrow\sqrt{T_{k}(t-1)}\hat{\Delta}_{k,\epsilon}(t-1)$
@@
Let $K>0$ and $T\geq 2K$. Assume that all the arms have rewards in [0,1][0,1][ 0 , 1 ] with mean $\mu_{k}$. Let $\tau\in\mathbb{R},\zeta\geq 0,\epsilon\geq 0$. The expected loss of Algorithm 1 is upper bounded by
@@ 4.1 Algorithm for the Fixed Confidence Case
6:  Compute confidence radius $\text{rad}(i)\,\,\forall i\in[K]$triangleright where $\text{rad}_{t}(i)=\sqrt{\frac{\log\left(\frac{4Kt^{3}}{\delta}\right)}{8T_{i}(t)}}$
@@
Let $K>0$, and let all the arms rewards be in [0,1][0,1][ 0 , 1 ]. Let $\tau\in\mathbb{R}$, $\epsilon\geq 0$, $\delta\in(0,1)$. Then, with probability at least $1-\delta$, algorithm 2 returns the optimal set and
@@ 5 Conclusion
Let arm $a$ with mean reward $\mu_{a}$ and privacy parameter $\epsilon>0$ be given. Let $R$ denote the reward returned by pulling the arm $a$ and use $B(R)$ to denote the output of $PrivBern(\epsilon)$ observed by the agent. The value of $B$ is either 111 or 0.
@@
where $\frac{\sqrt{2}\delta}{1-2\sqrt{2\delta}}$ simplifies to 1/21/21 / 2 for $\delta=(4\sqrt{2})^{-1}$.
@@
  • $[\hat{\mu}_{i}^{t}-\text{rad}_{t}(i),\hat{\mu}_{i}^{t}]$ if $i\in S_{t}$, i.e., if the empirical mean is above the threshold,
  • $[\hat{\mu}_{i}^{t},\hat{\mu}_{i}^{t}+\text{rad}_{t}(i)]$ if $i\notin S_{t}$, i.e., if the empirical mean is below the threshold.

### v2

@@ 2.2 Thresholding Bandits
In the fixed budget thresholding bandit problem (Locatelli et al., 2016), the goal is to identify a subset of arms from a set $\mathcal{A}=\{1,2,\dots,K\}$ whose mean rewards exceed a prespecified threshold, up to a tolerance factor $\zeta$. Formally, let each arm $i$ in $\mathcal{A}$ have a reward that is distributed within $[0,1]$ and an unknown mean reward $\mu_{i}$. Let $\tau$ denote the desired threshold, and let $\Delta_{i}=|\tau-\mu_{i}|+\zeta$ represent the deviation of the mean reward of arm $i$ from the threshold plus some tolerance $\zeta\geq 0$. The objective is to return $\mathcal{S}_{\tau}\subseteq\mathcal{A}$ where:
@@ 2.3.1 Bernoulli Mechanism
The Bernoulli mechanism, which will be utilized in both the fixed budget (Section 3.1) and fixed confidence (Section 4.1) settings, transforms a bounded reward in $[0,1]$ into a Bernoulli random variable that satisfies differential privacy guarantees. This mechanism produces a binary output (0 or 1) that serves as a privacy-preserving approximation of the reward. Consequently, the agent does not have direct access to the true reward of the arm; instead, they observe only the realization of a Bernoulli random variable.
@@
Given a reward $r\in[0,1]$ to be privatized, the PrivBern($\epsilon$) Mechanism returns a binary value sampled from a Bernoulli distribution. Specifically, the mechanism:
@@
The PrivBern($\epsilon$) Mechanism is an $\epsilon$-DP mechanism that takes as input a reward in $[0,1]$ from arm $a$, and outputs a Bernoulli-distributed random variable with an expected value
@@ 3.1 Algorithm for the Fixed Budget Case
where $T_{k}(t-1)$ is the number of pulls of arm $k$ up to time t1t-1 and
@@
is the estimate of arm $k$’s private reward deviation from $\tau_{\epsilon}$ at time t1t-1, and
@@
6:  Pull arm $I_{t}\leftarrow\arg\min_{k\leq K}B_{k}(t)$ $\triangleright$ where $B_{k}(t)\leftarrow\sqrt{T_{k}(t-1)}\hat{\Delta}_{k,\epsilon}(t-1)$
@@
Let $K>0$ and $T\geq 2K$. Assume that all the arms have rewards in $[0,1]$ with mean $\mu_{k}$. Let $\tau\in\mathbb{R},\zeta\geq 0,\epsilon\geq 0$. The expected loss of Algorithm 1 is upper bounded by
@@ 4.1 Algorithm for the Fixed Confidence Case
6:  Compute confidence radius $\text{rad}(i)\,\,\forall i\in[K]$ $\triangleright$ where $\text{rad}_{t}(i)=\sqrt{\frac{\log\left(\frac{4Kt^{3}}{\delta}\right)}{8T_{i}(t)}}$
@@
Let $K>0$, and let all the arms rewards be in $[0,1]$. Let $\tau\in\mathbb{R}$, $\epsilon\geq 0$, $\delta\in(0,1)$. Then, with probability at least $1-\delta$, algorithm 2 returns the optimal set and
@@ 5 Conclusion
Acknowledgments
Annalisa Barbara acknowledges support from the EPSRC Centre for Doctoral Training in Statistics and Machine Learning EP/Y034813/1 and from the MUR - PRIN 2022 project 2022R45NBB funded by the Next Generation EU program.
@@ Acknowledgments
Let arm $a$ with mean reward $\mu_{a}$ and privacy parameter $\epsilon>0$ be given. Let $R$ denote the reward returned by pulling the arm $a$ and use $B(R)$ to denote the output of $PrivBern(\epsilon)$ observed by the agent. The value of $B$ is either $1$ or 0.
@@
where $\frac{\sqrt{2}\delta}{1-2\sqrt{2\delta}}$ simplifies to $1/2$ for $\delta=(4\sqrt{2})^{-1}$.
@@
  • $\hat{\mu}_{i}^{t}-\text{rad}_{t}(i),\hat{\mu}_{i}^{t}]$ if $i\in S_{t}$, i.e., if the empirical mean is above the threshold,
  • $\hat{\mu}_{i}^{t},\hat{\mu}_{i}^{t}+\text{rad}_{t}(i)]$ if $i\notin S_{t}$, i.e., if the empirical mean is below the threshold.